# Do Generative AI Models Output Harm while Representing Non-Western Cultures: Evidence from A Community-Centered Approach

### v2

@@ 5.2 Honoring Heterogeneity within Cultures
However, as our participants demonstrated, Indian-ness is not a singular identity: there exist various subcultures with distinct features and traditions grounded in centuries of historical context. Our findings are full of instances where such diversity is homogenized, such as Satyajit’s remarks that “[the output is] not capturing region-specific phenomena or region-specific features, [and] that’s a drawback.” Any efforts to depict Indian-ness within T2I model outputs must honor the heterogeneity and diversity of India. We align ourselves with Qadri et al. (2023) in calling for stronger community-centered approaches towards designing T2I tools, and similarly advocate for the involvement of epistemic experts in the design process. Indian people are experts on their Indian-ness, and community-centered practices of data collection and annotation are an important step towards reducing harm within the outputs of T2I tools (Gadiraju et al. 2023).
However, asking individuals to be experts or representatives of a culture/subculture also has its own flaws, the chief of which is the fact that this ask risks performing the same homogenization that this paper finds fault in. Our participants recognized this dichotomy, and offered other ways of honoring heterogeneity within Indian culture. One such approach was the proposed fine-tuning of models for specific cultures/subcultures, instead of advertising them for widespread use. Participants noted that in trying to access a larger market, GAI tools risked causing harm to a wide group of users and losing a customer base. Taking inspiration from GAI tools such as Lesan, a machine translation tool and chatbot for Ethiopian languages which promises stronger results than tools from the Google Suite or ChatGPT (Deck 2023), we advocate for stronger support and providing resources for designers of GAI tools for specific contexts, embedded within the culture they are trying to serve.
At the bare minimum, as models continue to homogenize depictions of Indian-ness, participants suggested that the models should acknowledge their limitations, recognize the diversity of cultures and subcultures, and provide disclaimers for inaccurate representations. Aravind mentioned, ‘‘It would have been good if it had a single line recognizing the fact that it’s not composed of all the different cultures or subcultures that it’s trying to describe.”
@@ 5.3 Accuracy and Cultural Sensitivity
In analyzing behaviours of sociotechnical GAI models representing Indian subcultures, participants identified a unique facet related to model behavior diverging from expected actions, particularly concerning the given prompts. These deviations affect the model’s accuracy within the lens of machine learning. The model creates erroneous responses based on the inputs, portraying them as factual wherein it was completely against what was asked for. An example of this phenomenon occurred when participants prompted the generation of images depicting Indian women separately in Western attire, modern wear, or non-traditional clothing. Despite these specific prompts, the model consistently generated images of women in traditional sarees, contrary to expectations. These examples highlight the need for GAI tools to accurately interpret and respond to prompts. Participants emphasized the need for AI models to generate contextually appropriate responses that align with the prompt, demonstrate multilingual understanding by moving beyond weak transliterations and acknowledging diverse linguistic and cultural contexts, and avoid generating harmful or wrongful content.
In response to the concerns and recommendations outlined, we have developed a comprehensive set of guidelines for creating culturally sensitive AI models, as shown in Figure 4. These guidelines, presented in detail in the Appendix B, provide a framework for developers and researchers to consider and adhere to, ensuring that their models prioritize cultural understanding, accuracy, and inclusivity. These considerations are rooted in our findings around Indian culture/subculture, but are structured to be adaptable for any cultural group’s use. The recommendations ensure that GAI systems are designed with a deeper understanding of the cultural contexts and nuances that shape user interactions.
@@ 6 Discussion
In this paper, we structure our narration around sociotechnical AI systems, exploring their impact on social and technical actors within specific ecosystems. These systems, exemplified by the widespread use of technical models like NLP and computer vision in social domains such as education and policy-making, often operate as opaque ‘black boxes’ (O’neil 2017). This lack of transparency means that many users need to be made aware of the underlying mechanisms and decision-making processes of these models, leading to instances of discrimination in social contexts. Recognizing this, our research was motivated by the need for a community-based study focused on generative AI models released and sociotechnical system for public consumption.
@@
Towards this end, our framework (Figure 4) proposes key considerations for developing culturally sensitive GAI tools. In spite of the fact that such models will likely be universally used and likely embedded within suites of AIaaS, we do not advocate for theories of universal design that seek to produce one-size-fits-all solutions. Instead, rooted in participant observations, we advocate for fine-grained tools for specific contexts, and focus on designing for non-Western cultures.
7 Limitations
@@
8 Conclusion
@@
Acknowledgement
This work was supported by the U.S. National Institute of Standards and Technology (NIST) Grant 60NANB23D194. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily replect those of NIST.
Appendix A Additional Images
Figure 5: Stable Diffusion output for the prompt ‘Indian people in Western attire’.
Figure 6: Stable Diffusion output for the prompt ‘Indian people in Suits’.
Figure 7: Stable Diffusion output for the prompt ‘Indian women in Suits’.
Figure 8: Stable Diffusion output for the prompt ‘Indian woman’.
Appendix B Community-Based Design Recommendations
@@
A. Valuing the Evolution of Cultures:
@@
B. Honoring Heterogeneity within Cultures
@@
C. Ensuring Accuracy and Cultural Sensitivity in Responses
@@
Appendix C Participant Details
No. Psuedonym Gender Subculture Focus

### v3

@@ 5.2 Honoring Heterogeneity within Cultures
However, as our participants demonstrated, Indian-ness is not a singular identity: there exist various subcultures with distinct features and traditions grounded in centuries of historical context. Our findings are full of instances where such diversity is homogenized, such as Satyajit’s remarks that “[the output is] not capturing region-specific phenomena or region-specific features, [and] that’s a drawback.” Any efforts to depict Indian-ness within T2I model outputs must honor the heterogeneity and diversity of India. We align ourselves with Qadri et al. (2023) in calling for stronger community-centered approaches towards designing T2I tools, and similarly advocate for the involvement of epistemic experts in the design process. Indian people are experts on their Indian-ness, and community-centered practices of data collection and annotation are an important step towards reducing harm within the outputs of T2I tools (Gadiraju et al. 2023; Mack et al. 2024).
However, asking individuals to be expert representatives of a culture/subculture has its own flaws, like the fact that this risks performing the same homogenization that this paper faults. Our participants recognized this dichotomy, and offered other ways of honoring heterogeneity within Indian culture. One such approach was the proposed fine-tuning of models for specific cultures/subcultures, instead of advertising them for widespread use. Participants noted that in trying to access a larger market, GAI tools risked causing harm to a wide group of users and losing a customer base. Taking inspiration from GAI tools such as Lesan, a machine translation tool and chatbot for Ethiopian languages which promises stronger results than tools from the Google Suite or ChatGPT (Deck 2023), we advocate for stronger support and providing resources for designers of GAI tools for specific contexts, embedded within the culture they are trying to serve. Participants suggested that the models should acknowledge their limitations beyond broad statements such as ‘our results may be biased,’ recognize the diversity of cultures and subcultures, provide disclaimers for inaccurate representations, and account for their impact. As Aravind effectively put it, ‘‘It would have been good if it had a single line recognizing the fact that it’s not composed of all the different cultures or subcultures that it’s trying to describe.”
@@ 5.3 Accuracy and Cultural Sensitivity
In analyzing behaviours of sociotechnical GAI models representing Indian subcultures, participants identified a unique facet related to model behavior diverging from expected actions, particularly concerning the given prompts. These deviations affect the model’s accuracy within the lens of machine learning. The model creates erroneous responses based on the inputs, portraying them as factual wherein it was completely against what was asked for. An example of this phenomenon occurred when participants prompted the generation of images depicting Indian women separately in Western attire, modern wear, or non-traditional clothing. Despite these specific prompts, the model consistently generated images of women in traditional sarees, contrary to expectations. These examples highlight the need for GAI tools to accurately interpret and respond to prompts. Participants also emphasized the need for AI models to generate contextually appropriate responses that align with the prompt, demonstrate multilingual understanding by moving beyond weak transliterations and acknowledging diverse linguistic and cultural contexts, and avoid generating harmful content.
In response to the concerns and harms outlined above, we provide a set of guidelines for creating culturally sensitive AI models, as shown in Figure 4. These guidelines, presented in detail in Appendix A.1, provide a framework for developers and researchers to consider and adhere to, ensuring that their models prioritize cultural understanding, accuracy, and inclusivity. These considerations are rooted in our findings around Indian culture/subculture, but are adaptable for any underrepresented cultural group’ towards more accurate representation of their contexts in GAI outputs.
@@ 6 Discussion
In this paper, we structure our narration around sociotechnical AI systems, exploring their impact on social and technical actors within specific ecosystems. These systems, exemplified by the widespread use of GAI tools in domains domains that impact society and people’s life outcomes and opportunitie, such as education and policy-making, often operate as opaque ‘black boxes’ (O’neil 2017). This lack of transparency means that many users need to be made aware of the underlying mechanisms and decision-making processes of these models, leading to instances of discrimination in social contexts. Recognizing this, our research was motivated by the need for a community-based study focused on GAI and sociotechnical systems for public consumption.
@@
Towards this end, our framework (Figure 4) proposes key considerations for developing culturally sensitive GAI tools. In spite of the fact that such models will likely be ubiquitous and embedded within suites of AIaaS, we do not advocate for theories of universal design that seek to produce one-size-fits-all solutions. Instead, rooted in participant observations, we advocate for fine-grained tools for specific contexts, and focus on designing for non-Western cultures.
7 Limitations, Future Work and Conclusion
@@
Furthermore, this study is conducted using English prompts asking for representations of Indian culture. Although English is widely spoken in India and a lot of GAI usage is done so in English, future iterations of this study can experiment with prompts in Indian languages.
@@
Acknowledgements
This work was supported by the U.S. National Institute of Standards and Technology (NIST) Grant 60NANB23D194. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect those of NIST.
We are also appreciative of our reviewers for their thoughtful feedback, which undoubtedly made this paper richer than it otherwise would have been.
Appendix A Appendix
In this section, we provide complementary resources and experimental information that were omitted from the primary narration of the paper. This includes additional visualizations and a detailed description of the publicly accessible repository containing our codebook and the comprehensive collection of images generated during the focus groups. This supplementary material is intended to provide further context and transparency, supporting the reproducibility and replicability of our research.
A.1 Community-Based Design Recommendations
@@
I. Valuing the Evolution of Cultures:
@@
II. Honoring Heterogeneity within Cultures
@@
III. Ensuring Accuracy and Cultural Sensitivity
in Responses
@@ III. Ensuring Accuracy and Cultural Sensitivity
A.2 Additional Images
A.3 Participant Details
Figure 5: Stable Diffusion output for the prompt
‘Indian people in Western attire’.
Figure 6: Stable Diffusion output for the prompt
‘Indian people in Suits’.
Figure 7: Stable Diffusion output for the prompt
‘Indian women in Suits’.
Figure 8: Stable Diffusion output for the prompt
‘Indian woman’.
Figure 9: Stable Diffusion output for the prompt
‘Indian God’.
No. Psuedonym Gender Region of India Focus Group No.